Points given for fast correct answers thank you!
ently sells for $39.00 per share. The growth rate is a constant 9%, and the Banyan Co.s common stock curr company has an expected dividend yield of 6%. The expected long-run dividend payout ratio is 25%, and the expected return on equity (ROE) is 12%. New stock can be sold to the public at the current price, but a flotation cost of 15% would be incurred, what would be the cost of new equity? Round your answer to two decimal places. Do not round your intermediate calculations.
Answer #1

Solution-

Current price (P0)= $39

Growth rate (g)=9%

Dividend yield=6%

Next Expected dividend (D1)= $39*6%= $2.34

Flotation cost (F) =15%

Cost of new equity = D1/P0(1-F) + g

Cost of new equity = 2.34/39(1-0.15) + 0.09= 16.06%